A manager for an insurance company believes that customers have the following preferences for life insurance products: 40 % p Find the test statistic and the critical value.
Answer #1

We use a goodness of fit test

H0: The sample has a distribution which agrees with the expected distribution.

Ha: The sample has a distribution that is different from the expected distribution.

The Test Statistic:

Observed Expected % Expected (O-E)2 (O-E)2/E
Whole 77 40 84.4 54.76 0.64880
Universal 77 30 63.3 187.69 2.96510
Annuities 57 30 63.3 39.69 0.62700
Total 211 100 211 4.24

\chi^2 Test = 4.24

The critical value for df = n - 1 = 3 - 1 = 2, for Alpha = 0.05 is 5.992

(If required for Alpha = 0.01, then critical value = 9.210)

__________________________________________

Since \chi^2 Test < \chi^2 Critical, We fail to reject H0.